OpenStudy (anonymous):

-3n-5+n=4n+13 showing all work

OpenStudy (anonymous):

-3n-5-3n-13=0 -6n-18=0 -6n=18 n=-3

OpenStudy (anonymous):

-3n-5-3n-13=0 -6n-18=0 -6n=18 n=-3<<<<same